# Path-integral Monte Carlo estimator for the dipole polarizability of quantum plasma

### v1

@@ II.2 Analytic continuation of the Drude model
$G^{D}_{\mu\mu}(i\omega_{n})=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right].$ (22)
The analytical expression given by Eq.  (22) is valid for $\omega_{n}>0$. The static limit is defined here based on the zero-frequency Fourier transform of relative correlation function
$\tilde{G}^{D}_{\mu\mu}(i\omega_{n})=\int_{0}^{\beta}d\tau\tilde{G}^{D}_{\mu\mu}(\tau),$ (23)
but as discussed before, it is not physical and only used for numerical assessment of performance.
From Eq. (20) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(5). To consider the interacting problem, let us therefore turn to numerical path-integral simulation.
@@ II.3 PIMC estimator
The PIMC method can be used to simulate interacting quantum particles at finite temperatures. The method draws samples from the distribution of the thermal density matrix:
@@
where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H}$ is the density operator at inverse temperature $\beta$, $H$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function. Equation (24) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles. The boltzmannons evade the FSP but do not correspond to physical electrons, alas, except at high temperatures and low densities [23] (or strong coupling [32]), where particle repulsion outweighs exchange and the quantum statistics merge. Our results are also affected by the (lack of proper) quantum statistics.
@@
$H(R)=T+V=\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (25)
where $\lambda_{m}=1/2m_{e}$ and $T$ and $V$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles. Since $T$ and $V$ do not commute, the density operator must be expanded with the semi-group property
$e^{-\beta(T+V)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau T}e^{-\Delta\tau V}\right],$ (26)
where $\Delta\tau=\beta/M$ is the time-step and $M$ is the Trotter number. Using Eq. (26) and the product property [23], the exact Coulomb action between pairs of species can be treated for finite time-steps [33]. In this so-called pair approximation, higher-body interactions retain a modest time-step error proportional to $(\Delta\tau)^{3}$ [34, 33]. Therefore, it is justified to choose finite $M$, making the PIMC integrand a discrete walker of $M$ times $R$ coordinates, which forms a closed loop, an imaginary-time trajectory:
@@
The calculation of the thermal expectation values of diagonal properties, such as potential energy $V$, is done from Eq. (27) as
$\langle V\rangle$ $=Z^{-1}\int dR\langle R|\hat{V}\hat{\rho}(\beta)|R\rangle$ (28)
$=Z^{-1}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)V(R_{0})$ (29)
$=\frac{1}{M}\sum_{j=1}^{M}\langle V(R_{j}),$ (30)
where the last line employs symmetry properties of the trace, enabling to calculate mean over all time-slices of any trajectory. Because of the kinetic operator, the total energy $E$ is more complicated; here we use the thermal estimator [23], whose numerically exact pair energy $\partial\rho/\partial\beta$ can be precomputed [33].
The calculation of time-ordered observables, such as the imaginary-time correlation function from Eq. (12), is equally straightforward. The mean per-particle correlation function is evaluated as
$\langle G_{\mu\mu}(\tau)\rangle=\frac{1}{ZN}\prod_{j=1}^{M}dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu(R_{0})\mu(R_{t}),$ (31)
where the time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$. The expression can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [28, 29]. In the path-integral context, the dipole moment operator is
$\mu(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left[\mu(\mathbf{r}_{i}(t\Delta\tau)-\langle\mu_{i}\rangle\right],$ (32)
which again depends on the point of reference $\langle\mu_{i}\rangle$. However, as was done in Ref. [26], we identify
$\langle\tilde{G}_{\mu\mu}(\tau)\rangle$ $=\frac{1}{2N}\langle(\mu(0)-\mu(\tau))^{2}\rangle,$ (33)
where the relative correlation function arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path.
@@ II.4 Periodic boundary conditions
The simulation of charged plasma requires periodic boundary conditions (PBC). Here we use a cubic box of size $L$, whose size is determined by
$L=r_{s}\left(\frac{4\pi N}{3}\right)^{(1/3)},$ (34)
where $r_{s}=(3/4\pi n)^{1/3}$ is the Wigner-Seitz radius of the particle gas. The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [36, 37]. Note that the long-wavelength approximation discussed in Sec. II.1 also uses $L$, and the symbols are in practice interchangeable but semantically different.
The finite density is enforced for the many-body interactions of the plasma, but otherwise the evaluation resembles an open-boundary problem. So, the evaluation of the dipole polarizability estimator, Eq. (33), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries. Since the simulated particles may not exchange paths (permute), non-zero winding of a path through the cell is an artifact of the finite size. Therefore, we reject all moves leading to non-zero winding. Beside this constraint, the sampling is done using the multilevel bisection method [23]. The probability of winding increases when the thermal wavelength of the bisection move approaches the cell dimension, i.e.
@@
where $l$ is the bisection multilevel. We constrain $l$ so that the l.h.s. of Eq. (35) is lower than $L/8$. This is a simplification to avoid using periodic free-particle density matrix [23, 33], which would not inhibit the winding, anyway. As a result, fewer than one winding move per $10^{3}$ trial moves is observed and has to be rejected in any density. A modest sampling bias is mitigated along other finite-size effects when $L$ naturally grows with $N^{1/3}$.
@@ III Computational details
The PIMC simulation software, including the exact Coulomb pair action generation, implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring parallel sampling and data binning of a number of Markovian walkers. Here, 40 CPU cores are used for parallel sampling. Raw data after sample binning is stored in hierarchical data format (HDF5) and data processing chains operate on standard numerical Python libraries. The integral transformations, such as Eq. (16), are carried out using numerical quadrature. The simulation campaigns are operated with Nexus workflows [38] on various high-performance computing (HPC) facilities listed in Sec. V. All implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [39].
@@ IV Results
Figure 1: Comparison overview of the relative dipole correlation function $\tilde{G}_{\mu\mu}$ from PIMC and the Drude model. In the top, the imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$. In the bottom, the first few corresponding Matsubara data $\tilde{G}_{\mu\mu}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve. In both cases, the data match the analytically continued Drude model within statistical uncertainties.
@@
Let us now discuss results of the PIMC simulation, laid out in the earlier sections, as applied to charged quantum plasma at various finite temperatures and densities. With $m=m_{e}$ and $q^{2}=e^{2}$, the simulated particles resemble free electrons, only without sign-flipping permutations. Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector. The covered densities range between $r_{s}=2.0\ldots 8$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
The main observables of interest are overviewed in Fig. 1: the relative imaginary-time correlation function $\tilde{G}_{\mu\mu}(\tau)$ and its Matsubara series $G_{\mu\mu}(i\omega_{n})$. As stated in Sec. II, the vertical offset of $G_{\mu\mu}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\tilde{G}_{\mu\mu}(\tau)<0$ for all $\tau$. Again, the static Matsubara frequency $\omega_{n}=0$ must necessarily come out negative:
$G_{\mu\mu}(i\omega_{n}=0)=\int_{0}^{\beta}d\tau\tilde{G}_{\mu\mu}(\tau)<0,$ (36)
which stands out from the the rest, $G_{\mu\mu}(i\omega_{n}>0)>0$. Therefore, for convenience we will treat it with the opposite sign in figures and analyses, namely $\tilde{G}_{\mu\mu}(i\omega_{n}=0)\equiv-G_{\mu\mu}(i\omega_{n}=0)$. According to Eq. (11), $G_{\mu\mu}$ is directly connected to the physical polarizability for finite frequencies, i.e., $\alpha(\omega\neq 0)$. However, we will not seek here the ill-posed analytic continuation from imaginary to real domain.
@@
Rather, for validation of the results for $\tilde{G}_{\mu\mu}$, we compare the PIMC data to the Drude model, given in Eq. (22), in the imaginary domain. Thus, we report the relative error $\Delta\tilde{G}_{\mu\mu}=(G_{\mu\mu}-G^{D}_{\mu\mu})/G^{D}_{\mu\mu}$ for both the corresponding autocorrelation and Matsubara functions. Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability. As will be seen, the correspondence is close and $\Delta\tilde{G}_{\mu\mu}\neq 0$ is rarely statistically significant. Therefore, for the first-order proxy of accuracy we will assume
$\eta\equiv\Delta\tilde{G}_{\mu\mu}(i\omega_{n}=0),$ (37)
@@
Figure 2: Scaling of per-particle kinetic energy $K/N$ (left) and potential energy $V/N$ (bottom) with finite NN and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$. The solid lines correspond to raw PIMC observables and the dotted lines are finite-size corrected values based on Eqs. (38) and (39). The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so the $E_{p}$ is off above the plot.
Let us considering systematic merits of accuracy for the PIMC simulation. Finite-size effects of the energies are studied by varying the number of particles $N$ in the simulation, while keeping the density fixed. Results for kinetic and potential energies are presented in Fig. 2 for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$. Shown are also finite-size corrections [40, 34] ($\xi=0$), which rectify the energies toward the thermodynamic limit (TDL), according to:
@@
$V_{p}$ $=\frac{V_{N}}{N}+\frac{\omega_{p}}{4N},$ (39)
@@
where the corrected energies $K_{p}$ and $V_{p}$ are further multiplied with $\mathrm{tanh(\omega_{p}\beta)}$. The correlation energy is getting saturated with $N$, which also influences the kinetic energy. Simulation with $N>16$, or a reliable extrapolation to the TDL, would be needed to resolve the statistically significant finite-size bias of the energies. Also, clearly the Ewald summation with $k_{c}r_{c}>10$ is required to eliminate statistically significant bias from treating periodic images.
Figure 3: Scaling of $\tilde{G}_{\mu\mu}(\tau)$ (left) and $\tilde{G}_{\mu\mu}(i\omega_{n})$ (right) with finite NN. The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted). The bottom panels shows the relative differences $\Delta\tilde{G}_{\mu\mu}(\tau)$ (solid lines) and $\Delta\tilde{G}_{\mu\mu}(i\omega_{n})$, including statistical uncertainties (dotted lines).

### v2

@@ II.2 Quantum response functions
$\varepsilon_{0}\chi(\omega)=\left.\frac{\partial^{2}}{\partial E_{z}^{2}}\frac{\delta U_{E}(\omega)}{V}\right|_{E_{z}\rightarrow 0},$ (11)
which leads, using $U_{E}(\mathbf{r},\omega)=\mu(\mathbf{r})\cdot\mathbf{E}(\omega)$ and perturbation theory, to the Kubo formula [32]
$\chi(\omega)$ $=\frac{n}{\varepsilon_{0}}\alpha(\omega)$ (12)
$\alpha(\omega)$ $=-\frac{i}{\hbar}\int_{0}^{\infty}\mathrm{d}te^{i\omega t}\langle[\mu_{z}(t),\mu_{z}(0)]\rangle/N,$ (13)
where $\alpha(\omega)$ is per-particle dipole polarizability, $n=N/V$ is the number density, the angle brackets evaluate a quantum statistical average over the volume $V$, and square brackets denote a commutator. The dipole moment operator in $z$ direction is generally
$\mu_{z}(\mathbf{r},t)=\sum_{i=1}^{N}q_{i}z_{i}\delta(\mathbf{r}-\mathbf{r_{i}}(t)),$ (14)
where $q_{i}$ and $\mathbf{r_{i}}(t)$ are the respective charge and position of the $i$th particle. If $|q_{i}|=e$ for all $i$, Eqs. (12) and (13) can be expressed with the Fourier-space density–density susceptibility $\chi_{nn}(\omega)$ like in Eq. (1) or the dynamic structure factor [11], as is typically done to treat shorter wavelengths.
Let us denote the dipole–dipole correlation function with
$\frac{i}{\hbar}\langle[\mu(t),\mu(0)]\rangle/N\equiv G(t)$ (15)
such that the retarded (causal) correlation function is
$G^{R}(\omega)=\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}\Theta(t)G(t)=-\alpha(\omega),$ (16)
whose spectral density function is [25]
$A(\omega)\equiv-2\mathrm{Im}[G^{R}(\omega)].$ (17)
The spectral density allows analytic continuation of $G^{R}$ to imaginary time $it\rightarrow\tau$ through [25]
$\mathcal{G}(\tau)=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A(\omega),$ (18)
where $\beta=1/k_{B}T$ is the inverse temperature. The imaginary-time correlation function $\mathcal{G}(\tau)$ has bosonic symmetry with $\mathcal{G}(\tau)=\mathcal{G}(\tau+n\beta)$ for any integer $n$, and similar to its counterpart $G^{R}(t)$, is evaluated with
$\mathcal{G}(\tau)=\frac{1}{\hbar}\langle\mathcal{T}\mu_{z}(0)\mu_{z}(\tau)\rangle/N,$ (19)
where $\mathcal{T}$ is the time-ordering operator. The dipole moment operators $\mu_{z}(\tau)$ are also assumed normal ordered, meaning $\langle\mu_{Z}(\tau)\rangle=0$. The Fourier transform of $\mathcal{G}(\tau)$ is the Matsubara series $\mathcal{G}(i\omega_{n})$, where $\omega_{n}=2\pi n/\hbar\beta$ for all integers, which can also be obtained with [25]
$\mathcal{G}(i\omega_{n})=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1}{i\omega_{n}-\omega}A(\omega).$ (20)
While the imaginary-domain functions $\mathcal{G}(\tau)$ and $\mathcal{G}(i\omega_{n})$ contain the same information as $G^{R}(\omega)$, the conversion from imaginary to real domain by inverting Eqs. (18) or (20) is a known ill-posed problem, which is accentuated by the unavoidable error bars present in quantum Monte Carlo methods [25]. The use of existing numerical methods of analytic continuation, such as Maximum Entropy [26, 27], will not be pursued here.
Figure 1: Schematic illustration of the periodic dipole moment estimator and sampling constraints. Blue circles represent Monte Carlo walkers of quantum particles, i.e., positions advancing in discrete steps through imaginary time and forming closed loops $z(0)=z(\beta)$ modulo $L$, where $z$ denotes the real coordinate in one dimension. (left) The mean dipole autocorrelation function is easily measured by summing up individual particles, here represented by $\mu_{z}(\tau)/qN$ for $N=2$ by the smaller green dots. Their mean squared fluctuation w.r.t. to the mean (green vertical line) induces the relative autocorrelation function $\mathcal{\tilde{G}}(\tau)$ (visualized by the width of a green distribution). (middle) Unfolding of the path. Since the estimator does not treat periodicity with a Fourier phase, paths crossing the simulation box boundary to the nearest image must be unfolded, i.e., moved within $L/2$ of the earlier bead, before measurement. (right) Winding restriction. MC sampling may cause the path to wind across box boundary (branch of red dots), when the thermal wavelength of a multilevel bisection move is a significant fraction of $L$. For distinguishable particles, this is clearly an artifact of the finite cell and also detrimental to the estimation of $\mathcal{\tilde{G}}(\tau)$, and thus, such moves are rejected.
@@ II.3 PIMC estimator
The quantum statistical averages of the dipole–dipole correlation function at finite temperatures can be obtained with the canonical PIMC method [24]. The method draws samples from the distribution of the thermal density matrix:
@@
where $R$ denotes a set of variable particle coordinates, $\hat{\rho}(\beta)=e^{-\beta H_{0}}$ is the density operator at inverse temperature $\beta$, $H_{0}$ is the Hamiltonian and $Z=\langle R|\hat{\rho}|R\rangle$ is the thermal partition function. Equation (21) overlooks summations arising from spin degrees of freedom and particle permutations required to treat proper quantum statistics of bosons or fermions; instead, the simulation treats so-called Boltzmannons, which comprise a Coulomb interacting gas of distinguishable quantum particles [33]. The Boltzmannons evade the FSP but do not correspond to physical electrons, alas, except at high temperatures and low densities [24] (or strong coupling [34]), where particle repulsion outweighs exchange and the quantum statistics merge. As elaborated in the SI, some of our results indeed start to deviate from the Fermion values.
@@
$H_{0}=K+U=\hbar^{2}\lambda_{m}\sum_{i}\nabla_{i}^{2}+\sum_{i<j}\frac{q_{i}q_{j}}{4\pi\varepsilon_{0}|\mathbf{r}_{i}-\mathbf{r}_{j}|}$ (22)
where $\lambda_{m}=1/2m_{e}$ and $K$ and $U$ correspond, respectively, to the summation over kinetic and potential operators, the last ones comprising exact Coulomb interaction between pairs of particles. Since $K$ and $U$ do not commute, the density operator must be expanded with the semi-group property
$e^{-\beta(K+U)}=\lim_{M\rightarrow\infty}\left[e^{-\Delta\tau K}e^{-\Delta\tau U}\right],$ (23)
where $\Delta\tau=\hbar\beta/M$ is a finite time-step and $M$ is the Trotter number. Using Eq. (23) and the product property [24], the exact Coulomb action between pairs of species can be treated for finite time-steps [35]. In this so-called pair approximation, higher-body interactions retain a modest time-step error proportional to $(\Delta\tau)^{3}$ [36, 35]. Therefore, it is justified to choose finite $M$, making the PIMC integrand a discrete walker of $M$ times $R$ coordinates, which forms a closed loop, an imaginary-time trajectory:
@@
Thermal expectation values of diagonal properties, such as potential energy $V$, is can be calculated from
$\langle U\rangle$ $=Z^{-1}\int dR\langle R|U\hat{\rho}(\beta)|R\rangle$ (25)
$\approx Z^{-1}\int\left[\prod_{j=1}^{M}dR_{j}\rho\left(R_{j-1},R_{j};\Delta\tau\right)\right]U(R_{0})$ (26)
$=\frac{1}{M}\sum_{j=1}^{M}\langle U(R_{j})\rangle,$ (27)
where the last line employs symmetry properties of the trace, enabling to calculate the mean over all time-slices of any trajectory. Because of the kinetic operator, the total energy $E$ is more complicated; here we use the thermal estimator [24], whose numerically exact pair energy $\partial\rho/\partial\beta$ can be precomputed [35].
The calculation of time-ordered observables such as Eq. (19) is equally straightforward. The mean per-particle correlation function can evaluated at discrete multiples of $\Delta\tau$ by [30]
$\langle\mathcal{G}(\tau)\rangle=\frac{1}{\hbar ZN}\prod_{j=1}^{M}\int dR_{j}\rho(R_{j-1},R_{j};\Delta\tau)\mu_{z}(R_{0})\mu_{z}(R_{t}),$ (28)
where the integer time-difference index is $t=j+\tau/\Delta\tau$, modulo $M$. A snapshot measurement of the correlation function from the trajectory is visualized in Fig. 1 (left). The expression Eq. (28) can be easily symmetrized, optimized for evaluation and generalized for different combinations of measurement operators, as detailed in [29, 30].
The imaginary-time dipole moment operating on a time-slice is
$\mu_{z}(R_{t})$ $=\sum_{i=1}^{N}q_{i}\left(z_{t}-z_{0}\right),$ (29)
where $z_{0}$ is a reference point for $z$. A difficulty arises for free plasmas, for which $\sum_{i}q_{i}\neq 0$ and there is no extrinsic reference point. One might be tempted to use an intrinsic reference point, such as the path-centroid $z_{0}\equiv M^{-1}\sum_{t}z_{t}$, but that leads to $\langle\mathcal{G}(i\omega_{n}=0)=0$. The lack of a reference point is a fundamental consequence of the translation invariance, gracefully treated in the reciprocal space.
Therefore, we shall not directly evaluate $\langle\mathcal{G}(\tau)\rangle$ and instead (like in Ref. [27]) consider the relative correlation function
$\mathcal{\tilde{G}}(\tau)$ $\equiv\mathcal{G}(\tau)-\mathcal{G}(0)$ (30)
$=-\frac{1}{2\hbar}\langle(\mu(0)-\mu(\tau))^{2}\rangle/N,$ (31)
which arises from mean squared fluctuations of the instantaneous dipole moment along the imaginary-time path (the green path in Fig. 1 (left)). This is not such a huge compromise, as the relative correlation functions in real domain follow
$\tilde{G}^{R}(t)$ $\equiv G^{R}(t)-G^{R}(0)$ (32)
$\tilde{G}^{R}(\omega)$ $=\int dt\mathrm{e}^{i\omega t}\tilde{G}^{R}(t)=-\alpha(\omega\neq 0),$ (33)
so we can, in principle, treat all values of $\alpha(\omega)$ except the static value.
@@ II.4 Periodic boundary conditions
The simulation of charged plasma requires periodic boundary conditions (PBC). Here we use a cubic box of size $L$ (in practice like in Sec. II.1 although semantically different), whose size is determined by the
$L=r_{s}\left(\frac{4\pi N}{3}\right)^{1/3},$ (34)
where $r_{s}$ is the Wigner-Seitz radius. The treatment of periodic Coulomb images is treated using the ordinary Ewald summation [37, 38] (see SI for details). Evaluation of the relative dipole autocorrelation function (Eqs. (28) and (30)), is done on walkers whose all particle paths are unfolded into continuous trajectories near boundaries, as illustrated in Fig. 1 (middle). Since the simulated particles may not exchange paths (permute), non-zero winding of a path through the cell is an artifact of the finite size. Therefore, we reject all moves leading to non-zero winding (see Fig. 1 (right)). Beside this constraint, the sampling is done using the multilevel bisection method [24]. The probability of winding increases when the thermal de Broglie wavelength of the bisection move approaches the cell dimension, i.e.
@@
where $l$ is the bisection multilevel. We constrain $l$ so that the l.h.s. of Eq. (35) is lower than $L/8$. This is a simplification to avoid using periodic free-particle density matrix [24, 35], which would not inhibit the winding, anyway. As a result, fewer than one winding move per $10^{3}$ trial moves needs to be rejected in any density. A modest sampling bias is mitigated along other finite-size effects when $L$ naturally grows with $N^{1/3}$.
II.5 Analytic continuation of the Drude model
The ubiquitous model for considering the optical response free-carrier species like electrons in metals is the Drude model, whose dielectric susceptibility is
$\varepsilon^{D}(\omega)=1-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (36)
and thus, the susceptibility is
$\chi^{D}(\omega)=-\frac{\omega_{p}^{2}}{\omega^{2}+i\omega\Gamma},$ (37)
where $\omega_{p}^{2}=\tfrac{ne^{2}}{m\varepsilon_{0}}=\tfrac{e^{2}}{4\pi\varepsilon_{0}m}\tfrac{3}{r_{s}^{3}}$ is the plasma frequency and $\Gamma>0$ is a positive damping coefficient. As is well known [6] and iterated in Appendix A, the Drude model is recovered from the Lindhard function $\chi_{0}(\mathbf{q},\omega)$ in the limits of small $\Gamma$ and long wavelength as $q\rightarrow 0$. Thus, the Drude model is used in this work as a reference model that derives from quantum mechanics and contains screening on mean-field level [11] but neglects many-body effects from exchange and correlation. Differences between the PIMC data and Drude model are therefore manifestations of the many-body effects.
Based on Eqs. (12) and (37), the Drude polarizability per particle is given by
$\alpha^{D}(\omega)=\frac{\chi^{D}(\omega)\varepsilon_{0}}{n},$ (38)
whose spectral density, according to Eq (17), is
$A^{D}(\omega)$ $=\frac{2\omega_{p}^{2}\varepsilon_{0}}{n}\mathrm{Im}[\chi^{D}(\omega)]$ (39)
$=\frac{2e^{2}}{m}\frac{1}{\omega(\Gamma+\omega^{2}/\Gamma)}.$ (40)
From Eq. (39) it is clear that the Drude spectrum does not depend on the physical density, because the oscillators are not interacting; rather, the density only shows in the transformation to the macroscopic susceptibility with Eq.(12). The spectrum also diverges at $\omega=0$ complicating its analytic continuation into $\mathcal{G}$. However, the relative correlation function $\mathcal{\tilde{G}}$ can be obtained, because the subtraction of $G^{R}(\tau=0)$ controls the divergence at $\omega=0$. So, based on Eqs.(39), (18) and (20) we can perform numerical analytic continuation with
$\mathcal{\tilde{G}}^{D}(\tau)$ $=-\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\frac{1-\mathrm{e}^{-\tau\omega}}{1-\mathrm{e}^{-\beta\omega}}A^{D}(\omega).$ (41)
For the Matsubara data we get an analytic expression for the non-zero Matsubara frequencies:
$\mathcal{\tilde{G}}^{D}(i\omega_{n}\neq 0)$ $=\mathcal{G}^{D}(i\omega_{n}\neq 0)$ (42)
$=\frac{2\pi e^{2}}{m}\left[\frac{1}{\omega_{n}^{2}/\Gamma+\omega_{n}}\right],$ (43)
However, we can formally define the static limit based on the zero-frequency Fourier transform
$\mathcal{\tilde{G}}(i\omega_{n}=0)\equiv-\int_{0}^{\beta}d\tau\mathcal{\tilde{G}}(\tau),$ (44)
where the sign has been changed for convenience, to keep all values of $\mathcal{\tilde{G}}(i\omega_{n})>0$ in figures. As discussed before, $\mathcal{\tilde{G}}(i\omega_{n}=0)$ is not physical but since it describes the first-order scaling of the correlation function, we will use it later as a measure of quantity.
@@ III Computational details
The PIMC simulation software implements the canonical Metropolis Monte Carlo algorithm for a PIMC walker in Fortran90, featuring matrix squaring of the exact Coulomb pair action, parallel sampling and data binning of a number of Markovian walkers. Here, 40 CPU cores are used for the parallel sampling. Raw data after sample binning is stored in hierarchical data format (HDF5) and data processing chains operate on standard numerical Python libraries. The integral transformations, such as Eq. (18) and Eq. (20), are carried out using numerical quadrature. Fourier transforms from $\mathcal{\tilde{G}}(\tau)\rightarrow\mathcal{G}(i\omega_{n})$ are carried on regular grids of finite spacing using fast Fourier transform (FFT); however, FFT biases the result toward higher frequencies, and thus, the original data is interpolated 16-fold while truncating the size of the result array to mitigate the bias [26, 29]. The simulation campaigns are operated with Nexus workflows [39] on various high-performance computing facilities listed in Sec. VI. Central implementation details, including code snapshots and workflow scripts for step-by-step reproduction of the simulations, postprocessing and individual figures are supplied in a separate data repository [40].
@@ IV Results
The PIMC simulation, as laid out in earlier section, is applied in quantum plasmas at various finite temperatures and densities. With $m=m_{e}=1$ and $q^{2}=e^{2}=1$, the simulated particles resemble free electrons in the boltzmannon approximation. Thus, we use here conventional units of the HEG and express densities with $r_{s}$ and finite temperatures with $\Theta=T/T_{f}$, where $k_{B}T_{f}=\hbar^{2}k_{f}^{2}/2$ and $k_{f}=(3n\pi^{2})^{1/3}$ are, respectively, the Fermi temperature and the Fermi wavevector. The covered densities range between $r_{s}=2.0\ldots 12$ a0 and finite temperatures between $\Theta=T/T_{f}=0.1\ldots 0.8$.
The main observables of interest are overviewed in Fig. 2: the relative dipole autocorrelation function $\mathcal{\tilde{G}}(\tau)$ and its Matsubara series $\mathcal{\tilde{G}}(i\omega_{n})$. As argued in Sec. II, the vertical offset of $\mathcal{\tilde{G}}(\tau)$ is arbitrary, and thus, we treat it as zero, making $\mathcal{\tilde{G}}(\tau)\leq 0$ for all $\tau$. Thus, data for the static Matsubara frequency would be numerically negative unless changed defined with an opposite sign like in Eq. (44). In the absence of notable many-body effects, all values of $\mathcal{\tilde{G}}(i\omega_{n})$ lie on the same curve, except $\mathcal{\tilde{G}}(i\omega_{n}=0)$ which informs the first-order scale of the correlation function.
Figure 2: Comparison overview of the relative dipole correlation function $\mathcal{\tilde{G}}$ from PIMC and the Drude model. In the top, the imaginary-time correlation function $\mathcal{\tilde{G}}(\tau)$ is plotted at two temperatures, $\Theta_{1}=0.1$ and $\Theta_{2}=0.2$, resulting in two curves ending at different values for $\beta_{1}=2\beta_{2}$. In the bottom, the first few corresponding Matsubara data $\mathcal{\tilde{G}}(i\omega_{n})$ are plotted, where the frequencies $\omega_{n}=2\pi n/\hbar\beta$ for $\Theta_{1}$ are twice those of $\Theta_{2}$, but the two functions lie on the same curve. In both cases, the data match the analytically continued Drude model within statistical uncertainties.
Figure 3: Scaling of per-particle kinetic energy $K$ (left) and potential energy $V$ (bottom) with finite $N$ and variable dimensional Ewald cutoff parameter $r_{c}k_{c}$. The solid lines correspond to raw PIMC observables and the dotted lines are finite-size corrected values based on Eqs. (47) and (48). The case $r_{c}k_{c}=0$ means no analytical Ewald summation of any kind, including neutralizing background, and so $V$ is off above the plot.
We generally seek traces of methodological and physical discrepancies between the PIMC data and the Drude model. The differences are treated in terms of relative error
$\Delta\mathcal{\tilde{G}}=(\mathcal{\tilde{G}}-\mathcal{\tilde{G}}^{D})/\mathcal{\tilde{G}}^{D}$ (45)
@@
for any imaginary arguments $\tau$ or $i\omega_{n}$. Unless stated otherwise, we use $\Gamma=0.0001$ in the Drude model to approximate low-$\Gamma$ limit while maintaining numerical stability. As will be seen, the correspondence is close and $\Delta\mathcal{\tilde{G}}\neq 0$ is rarely statistically significant. Therefore, for a first-order measure of accuracy we shall use
$\eta\equiv\Delta\mathcal{\tilde{G}}(i\omega_{n}=0),$ (46)
@@
Figure 4: Scaling of $\mathcal{\tilde{G}}(\tau)$ (left) and $\mathcal{\tilde{G}}(i\omega_{n})$ (right) with finite $N$. The top panels show absolute values of the PIMC data (solid) and the corresponding Drude value (dotted). The bottom panels shows the relative differences $\Delta\mathcal{\tilde{G}}(\tau)$ (solid lines) and $\Delta\mathcal{\tilde{G}}(i\omega_{n})$, including statistical uncertainties (dotted lines).
However, let us first consider systematic merits of accuracy for the PIMC simulation based on energies. Finite-size effects of the energies are studied by varying the number of particles $N$ in the simulation, while keeping the density fixed. Results for kinetic and potential energies are presented in Fig. 3 and for up to $N=16$ and $\Theta=0.2$ at $r_{s}=4$. Shown are also finite-size corrections [41, 36] ($\xi=0$), which rectify the energies toward the thermodynamic limit (TDL), according to:
@@
$U_{p}$ $=\frac{U_{N}}{N}+\frac{\omega_{p}}{4N},$ (48)
@@
where the corrected energies $K_{p}$ and $U_{p}$ are further multiplied with $\mathrm{tanh}(\omega_{p}\beta)$. The correlation energy is getting saturated with $N$, which also influences the kinetic energy. Simulation with $N>16$, or a reliable extrapolation to the TDL, would be needed to resolve the statistically significant finite-size bias of the energies. Also, clearly the Ewald summation with $r_{c}k_{c}>10$ is required to eliminate statistically significant bias from treating periodic images. Details of the Ewald parameters and a table of the energy values is found in the SI.